OpenStudy (anonymous):

Identify the seventh term in the sequence in which a1=4096 and r=1/4

OpenStudy (anonymous):

\[a_n = a_1r^{n-1}\] so n = 7 (7th term), a_1 = 4096, and r = 1/4 gives: \[a_7=(4096)(\frac{1}{4})^{6}\] i dont have a calculator on me, one sec, while i wolfram that >.>

OpenStudy (anonymous):

well, it turns out the answer is 1 lol

OpenStudy (anonymous):

Thank you very much!